OpenStudy (anonymous):

Distribute on each side of the equation. 2. Collect like terms on each side of the equation. 3. Get all the variables on one side of the equation by doing the opposite math operation. 4. Solve for the variable. 8x - 1 = -4x + 23 what does x =?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

is there any bracket ????

OpenStudy (anonymous):

what do you mean

Nnesha (nnesha):

ok so you have to move x values on one side

OpenStudy (godlovesme):

8x+4x-1=-4x+4x+23 12x-1+1=23+1 12x=24 can u solve from there? @prowrestler

Nnesha (nnesha):

add 4x both side

Nnesha (nnesha):

 let me know what you get